Kyky232:

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 7 marbles, each of diameter 3 inches. The vase is filled with water up to a height of 12 inches. Which of the following could be used to calculate the volume of water in the vase? π(12in)2(3in) − 7(four over threeπ(1.5in)3) π(3in)2(12in) − 7(four over threeπ(1.5in)3) π(12in)2(3in) − 1.5(four over threeπ(7in)3) π(3in)2(12in) − 1.5(four over threeπ(7in)3)

Kyky232:

@snowflake0531

snowflake0531:

\[V = 2\pi r(h)\] And so we want to subtract the volume of where there's no water from teh volume of the whole vase

snowflake0531:

Wrote wrong, you have to subtract those weird marbles oops <.<

snowflake0531:

So we have the sphere volume as \[\frac{ 4 }{ 3 } \pi r^3\] and we have 7 of those so \[7(\frac{ 4 }{ 3 } \pi r^3)\]

snowflake0531:

So that is what we have to subtract

snowflake0531:

@snowflake0531 wrote:
\[V = 2\pi r(h)\] And so we want to subtract the volume of where there's no water from teh volume of the whole vase
pi r^2 i mean <.<

snowflake0531:

So we have to find the volume of hte whole cylinder, and subtract and 7 spheres

snowflake0531:

\[\pi r^2(h) - 7(\frac{ 4 }{ 3 } \pi r^3)\]

snowflake0531:

So now we put the h and r values in For the first one Height: 12 inches, diameter is 6, which means that radius is 3 \[\pi(3^2)(12)-7(\frac{ 4 }{ 3 }\pi r^3)\]

snowflake0531:

Then for the second one, we have diameter of 3, radius of 1.5 \[\pi(3^2)(12)-7(\frac{ 4 }{ 3 }\pi 1.5^3)\]

snowflake0531:

So, I think you got the answer now Ithink the choices were copied and pasted weirdly